OpenStudy (anonymous):

find the equation of the line passing through the point (-2, 4) and perpendicular to y = -x + 4.

OpenStudy (lalaly):

\[y-4=x-(-2)\]\[y=x+6\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

it didn't show up but it's not -x it's -1/2x

OpenStudy (lalaly):

-1/2? no itsnot

OpenStudy (anonymous):

like -one half x, I but in the fraction 1/2x but it only came up as -x

OpenStudy (lalaly):

oh lol

OpenStudy (anonymous):

what I mean is the question is y = -1/2x + 4

OpenStudy (anonymous):

yeah sorry!

OpenStudy (lalaly):

then the answer is\[y-4=2(x--2)\]\[y-4=2x+4\]\[y=2x+8\]

OpenStudy (anonymous):

isnt it suppose to be y=-2x+8?

OpenStudy (lalaly):

no coz the slope of the perpindicular line is the negative reciprocal and here its -1/2 so the negative reciprocal of -1/2 = 2

OpenStudy (anonymous):

ya, my fault, it suppose to be -(1/m).

OpenStudy (lalaly):

:)

OpenStudy (anonymous):

thanks guyss! (:

OpenStudy (anonymous):

got any idea how to solve this? \[\int\limits_{0}^{1/3} \tan ^{5} \pi \theta \sec \pi \theta d \theta\]